# Measurement of the branching fraction of $\eta\rightarrow\mu^{+}\mu^{-}$ and search for $\eta\rightarrow e^{+}e^{-}$

### v1

@@ I Introduction
Within the Standard Model (SM), the decay of $\eta$ into lepton pairs ($\eta\to\ell^{+}\ell^{-}$ with $\ell=e$ or $\mu$) proceeds via a fourth-order electromagnetic transition. The process is dominated by a two-photon intermediate state, as depicted in Fig. 1, resulting in significant suppression of the decay. Especially for $\eta\rightarrow e^{+}e^{-}$, the decay probability is expected to be very low, since the helicity factor is proportional to the square of the electron mass. Hence, a measurement of this decay probability higher than the SM model prediction could be indicative of hypothetical interactions arising from physics beyond the SM. This has stimulated considerable theoretical effort Young:1967zzb; BABU1982449; AMETLLER1983301; PhysRevD.29.911; PhysRevD.42.1509; gengLeptonicCPViolation1990; Margolis:1992gf; SAVAGE1992481; PhysRevLett.80.4633; Ametller:1993we; Knecht:1999gb; Silagadze:2006rt; Dorokhov:2009xs; Petri:2010ea devoted to calculating the decay rate of $\eta\to\ell^{+}\ell^{-}$. The predicted branching fractions for $\eta\to\mu^{+}\mu^{-}$ and $\eta\to e^{+}e^{-}$, based on the Canterbury approximants method Masjuan_2016, are $(4.72^{+0.05}_{-0.21})\times 10^{-6}$ and $(5.31^{+0.14}_{-0.04})\times 10^{-9}$, respectively.
Experimentally, a stringent upper limit of $\mathcal{B}(\eta\rightarrow e^{+}e^{-})<7\times 10^{-7}$, at 90% confidence level (CL), was obtained from the inverse process $e^{+}e^{-}\to\eta$ SND:2018egu. The world average of $\mathcal{B}(\eta\rightarrow\mu^{+}\mu^{-})=(5.7\pm 0.8)\times 10^{-6}$ PhysRevD.110.030001 is a combination of measurements performed by SATURNE II  PhysRevD.50.92 and Lepton-G DZHELYADIN1980471 more than thirty years ago. In recent years, taking advantage of the abundant production of $\eta,\eta^{\prime}$ mesons in $J/\psi$ radiative decays and the high statistics $J/\psi$ events, the BESIII experiment has made significant improvements in the study of their decays. However, the investigation of $\eta\to\ell^{+}\ell^{-}$ decays via the radiative decay $J/\psi\to\gamma\eta$ suffers from large backgrounds from $J/\psi\to\gamma\ell^{+}\ell^{-}$decays.
More recently, a novel approach Kang:2023yye was proposed to investigate $\eta$ decays, particularly for their rare or forbidden decays, with $\eta^{\prime}\rightarrow\pi\pi\eta$ produced in $J/\psi$ radiative decays. Therefore, using the sample of $(10087\pm 44)\times$ 106 BESIII:2021cxx $J/\psi$ events collected with the BESIII detector, we perform a study of $\eta\rightarrow\ell^{+}\ell^{-}$ with $J/\psi\rightarrow\gamma\eta^{\prime}$ and the subsequent decay of $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$.
@@ II BESIII DETECTOR and MONTE CARLO SIMULATION
The BESIII detector ABLIKIM2010345 is a magnetic spectrometer located at the Beijing Electron Positron Collider (BEPCII) ABLIKIM2010345. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The magnetic field was 0.9 T in 2012, which affects 11% of the total $J/\psi$ data. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identifier modules interleaved with steel. The charged-particle momentum resolution at 1 GeV/$c$ is 0.5%, and the specific ionization energy loss $\text{d}E/\text{d}x$ resolution is 6% for the electrons from Bhabha scattering. The EMC measures photon energies with a resolution of 2.5% (5%) at 1 GeV in the barrel (end-cap) region. The time resolution of the TOF barrel part is 68 ps, while that in the end-cap part is 110 ps. The end-cap TOF system was upgraded in 2015 using multi-gap resistive plate chamber technology, providing a time resolution of 60 ps Li:2017jpg; Guo:2017sjt; CAO2020163053, which benefits 87% of the data used in this paper.
Monte Carlo (MC) simulated data samples are produced with a geant4-based AGOSTINELLI2003250 software package, which includes the geometric description of the BESIII detector and the detector response. These samples are used to determine detection efficiencies and estimate the backgrounds. To thoroughly investigate potential backgrounds, we utilize an inclusive MC sample comprising 10 billion $J/\psi$ events. The inclusive MC sample includes both the production of the $J/\psi$ resonance and the continuum processes incorporated in kkmc JADACH2000260; PhysRevD.63.113009. All particle decays are modeled with evtgen LANGE2001152 using the branching fractions either taken from the Particle Data Group (PDG) PhysRevD.110.030001, when available, or otherwise estimated with lundcharm PhysRevD.62.034003. Final state radiation (FSR) from charged final state particles is incorporated using the photos package BARBERIO1991115. Specific generators using the appropriate theoretical amplitudes for $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\ell^{+}\ell^{-}$ Zhang:2012gt, $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ BESIII:2014bgm, and $\eta\rightarrow\gamma e^{+}e^{-}$ Qin:2017vkw decays were developed for this analysis.
@@ III EVENT SELECTION
For each candidate, particle identification (PID) is performed using the TOF and $\text{d}E/\text{d}x$ information, and four-constraint (4C) kinematic fits are performed imposing energy and momentum conservation under the hypotheses of $J/\psi\to\gamma\pi^{+}\pi^{-}\ell^{+}\ell^{-}$. The variable, $\chi^{2}_{\rm 4C+PID}=\chi^{2}_{4\rm C}+\chi^{2}_{\rm PID}$, is used to select the best candidate in events containing more than one combination. Here $\chi^{2}_{4\rm C}$ is the $\chi^{2}$ value of the 4C kinematic fit and $\chi^{2}_{\rm PID}$ is a sum over the four charged track candidates. For each event, the hypothesis with the smallest $\chi^{2}_{\rm 4C+PID}$ is kept for further analysis.
@@ IV Analysis of $\eta\rightarrow\mu^{+}\mu^{-}$
For the $\eta\rightarrow\mu^{+}\mu^{-}$ decay, $\chi^{2}_{\rm 4C+PID}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})<40$ is required. A 4C kinematic fit under the hypothesis of $J/\psi\rightarrow\gamma 2(\pi^{+}\pi^{-})$ is also performed, and $\chi^{2}_{\rm 4C+PID}(\pi^{+}\pi^{-}\pi^{+}\pi^{-})>\chi^{2}_{4\rm C+PID}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})$ is required to reject possible background events from $J/\psi\rightarrow\gamma 2(\pi^{+}\pi^{-})$. The muon counter information is not used due to the low momentum of the muons of interest.
@@
$\mathcal{B}(\eta\rightarrow\mu^{+}\mu^{-})$ $\,=\,\frac{N_{\rm obs}}{N_{J/\psi}\,{\mathcal{B}}_{\rm int}\,\varepsilon_{\eta\rightarrow\mu^{+}\mu^{-}}}$ (1)
@@
where $N_{\rm obs}$ is the signal yield, $N_{J/\psi}$ is the total number of $J/\psi$ events, $\mathcal{B}_{\rm int}\equiv\mathcal{B}(J/\psi\rightarrow\gamma\eta^{\prime})\,\mathcal{B}(\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta)$ is taken from the PDG PhysRevD.110.030001, $\varepsilon_{\eta\rightarrow\mu^{+}\mu^{-}}$ is the detection efficiency, and the uncertainty is statistical only.
@@ V Analysis of $\eta\rightarrow e^{+}e^{-}$
To select $\eta\to e^{+}e^{-}$ events, the $\chi^{2}_{4\rm C+PID}(\pi^{+}\pi^{-}e^{+}e^{-})$ is required to be less than 40. The $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass ($M_{\pi^{+}\pi^{-}e^{+}e^{-}}$) is shown in Fig. 5(a). No evident $\eta^{\prime}$ signal is observed. We then select the $\eta^{\prime}$ mass window to be within $(0.945,0.97)\,\text{GeV}/c^{2}$. Fig. 5(b) shows the $e^{+}e^{-}$ invariant mass ($M_{e^{+}e^{-}}$) distribution for candidates passing all selection criteria. No events are observed in the $\eta$ mass region.
@@
The upper limit on the signal yield, $N_{\rm UL}$, is estimated at a CL of 90% using the method from Ref. Rolke:2000ij, which is suitable for small signals in the presence of background. Given $N_{\rm sig}$ and $N_{\rm bkg}$, $N_{\rm UL}$ of $\eta\to e^{+}e^{-}$ is estimated to be 1.3. Accounting for efficiency and systematic uncertainty $\delta_{\rm sys}$ (see Section VI below for details), the upper limit on the branching fraction of $\eta\to e^{+}e^{-}$ is determined as
$\mathcal{B}(\eta\rightarrow e^{+}e^{-})$ $\,<\,\frac{N_{\rm UL}}{N_{J/\psi}\,\mathcal{B}_{\rm int}\,\varepsilon_{\eta\to e^{+}e^{-}}\,(1-\delta_{\rm sys})}$ (2)
@@ VI Systematic Uncertainties
    The total number of $J/\psi$ events, (10087 $\pm$ 44) $\times$ 106 BESIII:2021cxx, is determined from an analysis of inclusive $J/\psi$ hadronic events, and the uncertainty, 0.4%, is taken as the systematic uncertainty.
@@
    The branching fractions of $J/\psi\rightarrow\gamma\eta^{\prime}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$ are taken from the PDG PhysRevD.110.030001. Their uncertainties are 1.3% and 1.2%, respectively.
@@
    The single charged pion track efficiency has been studied with respect to the transverse momentum and polar angle $\theta$ in the MDC by means of a high-purity sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$ liuStudyTrackingEfficiency2025 obtained by using one pion to tag the presence of the other. Two-dimensional tracking efficiencies as a function of the transverse momentum and $\cos\theta$ are evaluated for both the data and MC simulation. To correct the MC efficiencies to those of the data, the corresponding two-dimensional correction factors and their uncertainties are derived. To estimate its impact on the systematic uncertainty, the MC simulation events are then weighted to determine the detection efficiency. Since the mass and the behavior of low-momentum muons are similar to those of pions, we weight the muon tracks by assuming them to be pions.
@@
    In case of $\eta\rightarrow\mu^{+}\mu^{-}$, the uncertainties from PID and the kinematic fit are evaluated separately. To estimate the pion PID efficiency uncertainty, we have studied the pion PID efficiency using the control samples of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$. The results show that the pion PID efficiency for data agrees well with that of the MC simulation. The minor MC-data discrepancy as a function of the momentum and the polar angle is used to estimate the systematic uncertainty with the same approach as for the tracking efficiency. Due to a lack of a specific control sample for low-momentum muons, the PID efficiency correction is performed with the correction factor obtained from pions. The uncertainty associated with the 4C kinematic fit is studied by correcting the track helix parameters PhysRevD.87.012002 to reduce the difference between data and MC simulation. The differences between the efficiencies with and without corrections, summarized in Table 1, are taken as the systematic uncertainties from PID and 4C kinematic fit.
    For $\eta\rightarrow e^{+}e^{-}$, we use a clean control sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0},\pi^{0}\rightarrow\gamma e^{+}e^{-}$ to study the uncertainty. This control sample has the same topology as the decay of interest. Using the same charged tracks and photon selection criteria, we first select the candidate events with a very loose requirement on $\chi^{2}_{4\rm C+PID}$. Then, we further tighten the selection by requiring $\chi^{2}_{4\rm C+PID}<40$. The efficiency difference between data and MC simulation, 4.6%, is taken as the systematic uncertainty.
@@
    In the fit to $M_{\mu^{+}\mu^{-}}$, the numbers of background events from $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ are fixed at the normalized values according to the branching fractions from the PDG PhysRevD.110.030001. To estimate the effect of the uncertainty of the used branching fractions, the numbers of events for these two backgrounds are varied randomly within the uncertainties of the branching fractions. Using these randomly scaled parameters, we perform a series of fits to the $M_{\mu+\mu^{-}}$ distribution. The difference of the determined number of signal events, 0.2%, is taken as the systematic uncertainty.

### v2

@@ I Introduction
Within the Standard Model (SM), the decay of $\eta$ into lepton pairs ($\eta\to\ell^{+}\ell^{-}$ with $\ell=e$ or $\mu$) proceeds via a fourth-order electromagnetic transition. The process is dominated by a two-photon intermediate state, as depicted in Fig. 1, resulting in significant suppression of the decay. Especially for $\eta\rightarrow e^{+}e^{-}$, the decay probability is expected to be very low, since the helicity factor is proportional to the square of the electron mass. Hence, a measurement of this decay probability higher than the SM model prediction could be indicative of hypothetical interactions arising from physics beyond the SM. This has stimulated considerable theoretical effort Young (1967); Babu and Ma (1982); Ametller et al. (1983); Scadron and Visinescu (1984); Geng and Ng (1990); Geng (1990); Margolis et al. (1993); Savage et al. (1992); Gómez Dumm and Pich (1998); Ametller et al. (1993); Knecht et al. (1999); Silagadze (2006); Dorokhov et al. (2009); Petri (2010) devoted to calculating the decay rate of $\eta\to\ell^{+}\ell^{-}$. The predicted branching fractions for $\eta\to\mu^{+}\mu^{-}$ and $\eta\to e^{+}e^{-}$, based on the Canterbury approximants method Masjuan and Sanchez-Puertas (2016), are $(4.72^{+0.05}_{-0.21})\times 10^{-6}$ and $(5.31^{+0.14}_{-0.04})\times 10^{-9}$, respectively.
Experimentally, a stringent upper limit of $\mathcal{B}(\eta\rightarrow e^{+}e^{-})<7\times 10^{-7}$, at 90% confidence level (CL), was obtained from the inverse process $e^{+}e^{-}\to\eta$ Achasov and others (2018). The world average of $\mathcal{B}(\eta\rightarrow\mu^{+}\mu^{-})=(5.7\pm 0.8)\times 10^{-6}$ Navas and others (2024) is a combination of measurements performed by SATURNE II  Abegg and others (1994) and Lepton-G Dzhelyadin and others (1980) more than thirty years ago. In recent years, taking advantage of the abundant production of $\eta,\eta^{\prime}$ mesons in $J/\psi$ radiative decays and the high statistics $J/\psi$ events, the BESIII experiment has made significant improvements in the study of their decays. However, the investigation of $\eta\to\ell^{+}\ell^{-}$ decays via the radiative decay $J/\psi\to\gamma\eta$ suffers from large backgrounds from $J/\psi\to\gamma\ell^{+}\ell^{-}$decays.
More recently, a novel approach Kang et al. (2023) was proposed to investigate $\eta$ decays, particularly for their rare or forbidden decays, with $\eta^{\prime}\rightarrow\pi\pi\eta$ produced in $J/\psi$ radiative decays. Therefore, using the sample of $(10087\pm 44)\times$ 106 Ablikim and others (2022) $J/\psi$ events collected with the BESIII detector, we perform a study of $\eta\rightarrow\ell^{+}\ell^{-}$ with $J/\psi\rightarrow\gamma\eta^{\prime}$ and the subsequent decay of $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$.
@@ II BESIII DETECTOR and MONTE CARLO SIMULATION
The BESIII detector Ablikim and others (2010) is a magnetic spectrometer located at the Beijing Electron Positron Collider (BEPCII) Ablikim and others (2010). The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The magnetic field was 0.9 T in 2012, which affects 11% of the total $J/\psi$ data. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identifier modules interleaved with steel. The charged-particle momentum resolution at 1 GeV/$c$ is 0.5%, and the specific ionization energy loss $\text{d}E/\text{d}x$ resolution is 6% for the electrons from Bhabha scattering. The EMC measures photon energies with a resolution of 2.5% (5%) at 1 GeV in the barrel (end-cap) region. The time resolution of the TOF barrel part is 68 ps, while that in the end-cap part is 110 ps. The end-cap TOF system was upgraded in 2015 using multi-gap resistive plate chamber technology, providing a time resolution of 60 ps Li and others (2017); Guo and others (2017); Cao and others (2020), which benefits 87% of the data used in this paper.
Monte Carlo (MC) simulated data samples are produced with a geant4-based Agostinelli and others (2003) software package, which includes the geometric description of the BESIII detector and the detector response. These samples are used to determine detection efficiencies and estimate the backgrounds. To thoroughly investigate potential backgrounds, we utilize an inclusive MC sample comprising 10 billion $J/\psi$ events. The inclusive MC sample includes both the production of the $J/\psi$ resonance and the continuum processes incorporated in kkmc Jadach et al. (2000, 2001). All particle decays are modeled with evtgen Lange (2001) using the branching fractions either taken from the Particle Data Group (PDG) Navas and others (2024), when available, or otherwise estimated with lundcharm Chen et al. (2000). Final state radiation (FSR) from charged final state particles is incorporated using the photos package Barberio et al. (1991). Specific generators using the appropriate theoretical amplitudes for $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\ell^{+}\ell^{-}$ Zhang et al. (2012), $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ Ablikim and others (2014), and $\eta\rightarrow\gamma e^{+}e^{-}$ Qin et al. (2018) decays were developed for this analysis.
@@ III EVENT SELECTION
For each candidate, particle identification (PID) is performed using the TOF and $\text{d}E/\text{d}x$ information, and four-constraint (4C) kinematic fits are performed imposing energy and momentum conservation under the hypotheses of $J/\psi\to\gamma\pi^{+}\pi^{-}\ell^{+}\ell^{-}$. For each event, the combination with the smallest $\chi_{\text{sum}}^{2}(\gamma\pi^{+}\pi^{-}\ell^{+}\ell^{-})$ is kept for further analysis, where $\chi_{\text{sum}}^{2}(\gamma\pi^{+}\pi^{-}\ell^{+}\ell^{-})=\chi_{\text{4C}}^{2}+\sum_{i=1}^{4}\chi_{\text{PID}(i)}^{2}$ is the sum of the 4C kinematic fit $\chi^{2}$ and PID $\chi^{2}$ of the four charged tracks.
@@ IV Analysis of $\eta\rightarrow\mu^{+}\mu^{-}$
For the $\eta\rightarrow\mu^{+}\mu^{-}$ decay, the events with $\chi^{2}_{\rm sum}(\gamma\pi^{+}\pi^{-}\mu^{+}\mu^{-})<40$ are retained. Here the requirement is optimized by the figure of merit Punzi $S/\sqrt{S+B}$, where $S$ is the number of signal events from MC simulation and $B$ is the number of background events in data. To reject possible background events from $J/\psi\rightarrow\gamma 2(\pi^{+}\pi^{-})$, $\chi^{2}_{\rm sum}(\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-})>\chi^{2}_{\rm sum}(\gamma\pi^{+}\pi^{-}\mu^{+}\mu^{-})$ is required. The muon counter information is not further used due to the low momentum of the muons of interest.
@@
$\mathcal{B}(\eta\rightarrow\mu^{+}\mu^{-})$ $\,=\,\frac{N_{\rm obs}}{N_{J/\psi}\cdot{\mathcal{B}}_{\rm int}\cdot\varepsilon_{\eta\rightarrow\mu^{+}\mu^{-}}}$ (1)
@@
where $N_{\rm obs}$ is the signal yield, $N_{J/\psi}$ is the total number of $J/\psi$ events, $\mathcal{B}_{\rm int}\equiv\mathcal{B}(J/\psi\rightarrow\gamma\eta^{\prime})\,\mathcal{B}(\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta)$ is taken from the PDG Navas and others (2024), $\varepsilon_{\eta\rightarrow\mu^{+}\mu^{-}}$ is the detection efficiency, and the uncertainty is statistical only.
@@ V Analysis of $\eta\rightarrow e^{+}e^{-}$
To select $\eta\to e^{+}e^{-}$ events, the $\chi^{2}_{\rm sum}(\gamma\pi^{+}\pi^{-}e^{+}e^{-})$ is required to be less than 40. The $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass ($M_{\pi^{+}\pi^{-}e^{+}e^{-}}$) is shown in Fig. 5(a). No evident $\eta^{\prime}$ signal is observed. We then select the $\eta^{\prime}$ mass window to be within $(0.945,0.97)\,\text{GeV}/c^{2}$. Fig. 5(b) shows the $e^{+}e^{-}$ invariant mass ($M_{e^{+}e^{-}}$) distribution for candidates passing all selection criteria. No events are observed in the $\eta$ mass region.
@@
The upper limit on the signal yield, $N_{\rm UL}$, is estimated at a CL of 90% using the method from Ref. Rolke and Lopez (2001), which is suitable for small signals in the presence of background. Given $N_{\rm sig}$ and $N_{\rm bkg}$, $N_{\rm UL}$ of $\eta\to e^{+}e^{-}$ is estimated to be 1.3. Accounting for efficiency and systematic uncertainty $\delta_{\rm sys}$ (see Section VI below for details), the upper limit on the branching fraction of $\eta\to e^{+}e^{-}$ is determined as
$\mathcal{B}(\eta\rightarrow e^{+}e^{-})$ $\,<\,\frac{N_{\rm UL}}{N_{J/\psi}\cdot\mathcal{B}_{\rm int}\cdot\varepsilon_{\eta\to e^{+}e^{-}}\cdot(1-\delta_{\rm sys})}$ (2)
@@ VI Systematic Uncertainties
    The total number of $J/\psi$ events, (10087 $\pm$ 44) $\times$ 106 Ablikim and others (2022), is determined from an analysis of inclusive $J/\psi$ hadronic events, and the uncertainty, 0.4%, is taken as the systematic uncertainty.
@@
    The branching fractions of $J/\psi\rightarrow\gamma\eta^{\prime}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta$ are taken from the PDG Navas and others (2024). Their uncertainties are 1.3% and 1.2%, respectively.
@@
    The single charged pion track efficiency has been studied with respect to the transverse momentum and polar angle $\theta$ in the MDC by means of a high-purity sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$ Liu and others (2025) obtained by using one pion to tag the presence of the other. Two-dimensional tracking efficiencies as a function of the transverse momentum and $\cos\theta$ are evaluated for both the data and MC simulation. To correct the MC efficiencies to those of the data, the corresponding two-dimensional correction factors and their uncertainties are derived. To estimate its impact on the systematic uncertainty, the MC simulation events are then weighted to determine the detection efficiency. Since the mass and the behavior of low-momentum muons are similar to those of pions, we weight the muon tracks by assuming them to be pions.
@@
    In case of $\eta\rightarrow\mu^{+}\mu^{-}$, the uncertainties from PID and the kinematic fit are evaluated separately. To estimate the pion PID efficiency uncertainty, we have studied the pion PID efficiency using the control samples of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$. The results show that the pion PID efficiency for data agrees well with that of the MC simulation. The minor MC-data discrepancy as a function of the momentum and the polar angle is used to estimate the systematic uncertainty with the same approach as for the tracking efficiency. Due to a lack of a specific control sample for low-momentum muons, the PID efficiency correction is performed with the correction factor obtained from pions. The uncertainty associated with the 4C kinematic fit is studied by correcting the track helix parameters Ablikim and others (2013) to reduce the difference between data and MC simulation. The differences between the efficiencies with and without corrections, summarized in Table 1, are taken as the systematic uncertainties from PID and 4C kinematic fit.
    For $\eta\rightarrow e^{+}e^{-}$, we use a clean control sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0},\pi^{0}\rightarrow\gamma e^{+}e^{-}$ to study the uncertainty. This control sample has the same topology as the decay of interest. Using the same charged tracks and photon selection criteria, we first select the candidate events with a very loose requirement on $\chi^{2}_{\rm sum}$. Then, we further tighten the selection by requiring $\chi^{2}_{\rm sum}<40$. The efficiency difference between data and MC simulation, 4.6%, is taken as the systematic uncertainty.
    Although the systematic uncertainty from PID and the 4C kinematic fit is larger than other individual contributions, its combination with the tracking efficiency correction yields a small net effect on the detection efficiency. We therefore use the result without the tracking efficiency correction as the nominal value and take the difference between the results with and without the correction as the systematic uncertainty.
@@
    In the fit to $M_{\mu^{+}\mu^{-}}$, the numbers of background events from $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ and $\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ are fixed at the normalized values according to the branching fractions from the PDG Navas and others (2024). To estimate the effect of the uncertainty of the used branching fractions, the numbers of events for these two backgrounds are varied randomly within the uncertainties of the branching fractions. Using these randomly scaled parameters, we perform a series of fits to the $M_{\mu+\mu^{-}}$ distribution. The difference of the determined number of signal events, 0.2%, is taken as the systematic uncertainty.